E-/D-Time

OSA: Steve Allanson

Isp: Jenny Power

Manager: Jeff Urban

ISA/Observer: B. Rothberg (Tucson)

AO: Doug Miller

========================================================

UT 01:13 - Opening

UT 01:14 - Switching LBC Software from Version 2.03 to 2.1

UT 01:30 - Doug is in 'da house

UT 02:00 - IT 5692 - No GCS images on FACSUM Big Board Display

UT 04:46 - completed MODS/PEPSI testing - moving to MODS1 for Rothberg D-time since LBC targets still too low

========================================================

UT 05:08 - Doing SpecStd first Feige 110

UT 05:17 - Seeing has exploded - wavering between 1.5-2, collimation is tough

UT 05:19 - Sievesnap had been taken and I reset data, so one file was overwritten (0001.fits) renamed the first blue to mods1b.20150925.0001b.fits

UT 05:31 - taking a 3rd specstd exposure (40degrees from Moon -also variable seeing)

UT 05:35 - Going to Science target FF0139+0115 -clouds on eastern horizon

UT 05:36 - Seeing ~ 1.5" @airmass 1.702

UT 05:49 - Taking spectra (can just barely see it in the slit as higher counts ~100-120+ over background at exactly Y=590)

UT 06:15 - first exposure can see emission lines and some continuum in Red, blue not so much - now at airmass 1.46

UT 06:16 - keeping an eye on the windspeeds -slowly rising, occasional 15m/s gusts

UT 06:29 - DIMM hitting 1.5-2"

UT 06:42 - Gusts getting to 20m/s - seeing still highly variable at airmass=1.35, DIMM 1.5-2.0, Guider also 1.5-2.0

UT 06:56 - taking a 4th exposure in blue and red airmass 1.3, seeing 1.3 DIMM 1.4 on Guider

UT 07:18 - taking a 5th exposure

UT 07:41 - moving to UGC 2238 - backup for backup MODS programs. seeing still highly variable 1.5-2.1" in GUIDER

UT 08:03 - DIMM 1.5"

UT 08:23 - Aborted second exposure Blue at 392sec, Red at 398sec - closing due to high wind gusts

UT 09:12 - Running MODS cals, pixel flats

UT 09:17 - Taking LBC Biases just in case we reopen (error on MODS pixel flats so start is really now)

UT 09:36 - Starting arc lamps

UT 09:45 - Starting slit flats

UT 10:06 - Starting BIASES 8kx3k

UT 12:00 - Turning off LBC and reverting back to LBC control 2.03

UT 12:04 - Putting MODS1 to sleep

-- %USERSIG{BarryRothberg - 2015-09-25}%

Comments

 
Topic revision: r4 - 25 Sep 2015, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback